### v3

@@ 7.1 General function approximation
To illustrate that the Squared ReLU activation in DSRN improves the derivative approximation, compared to neural networks using ReLU, we train DSRN and ReLU neural networks to learn the cubic function $f(x)=x^{3}$ as well as its first-order and second-order derivatives. In Figure 6, we visualize the cubic functions learned by the ReLU neural network (Figure 6(a)) and DSRN (Figure 6(d)), the first-order derivative $f^{\prime}(x)=3x^{2}$ estimated by the ReLU neural network (Figure 6(b)) and DSRN (Figure 6(e)), and the second-order derivative $f^{\prime\prime}(x)=6x$ estimated by the ReLU neural network (Figure 6(c)) and DSRN (Figure 6(f)). The comparison results demonstrate that the DSRN can approximate both the cubic function and its derivatives. In contrast, the ReLU neural network can only approximate the function up to the first-order derivative. Notably, the ReLU network fails to approximate the cubic function in terms of the second-order Sobolev norm, because its second-order derivative is zero, leading to significant errors in second-order approximation. On the other hand, the DSRN is able to learn the cubic function with the accurate first and second-order derivatives, achieving approximation in higher-order Sobolev norm.
(a) ReLU: $f(x)=x^{3}$
(b) ReLU: $f^{\prime}(x)=3x^{2}$
(c) ReLU: $f^{\prime\prime}(x)=6x$
(d) DSRN: $f(x)=x^{3}$
(e) DSRN: $f^{\prime}(x)=3x^{2}$
(f) DSRN: $f^{\prime\prime}(x)=6x$
@@ 7.2 One-dimensional Poisson equation
(81) $\Delta u(x)=-\pi^{2}\sin(\pi x),\quad x\in\left[-1,1\right],$
with boundary conditions $u(-1)=u(1)=0$. The exact solution is $u(x)=\sin(\pi x)$. To be specific, we approximate the solution $u(x)$ by the neural network $\mathcal{NN}_{\boldsymbol{\theta}}(x)$ and train the neural network parameters $\boldsymbol{\theta}$ to minimize the mean squared error(MSE)
$\frac{1}{N}\sum_{i=1}^{N}\left[\left(\frac{d^{2}}{dx^{2}}\mathcal{NN}_{\boldsymbol{\theta}}(x_{i})+\pi^{2}\sin(\pi x_{i})\right)^{2}+\left(\mathcal{NN}_{\boldsymbol{\theta}}(x_{i})-y_{i}\right)^{2}\right].$
@@
(a) ReLU
(b) DSRN
(c) Squared ReLU
(d) DSRN
@@
In Figure 7, We plot the prediction of the neural network with ReLU (Figure 7(a)), DSRN (Figure 7(b)) and the square of ReLU (Figure 7(c)), respectively. We also plot the first-order derivative $\frac{d}{dx}\mathcal{NN}_{\boldsymbol{\theta}}$ using the autograd. The exact solution is $u(x)=\sin(\pi x)$, which corresponds to the ground truth in the figure. Figure 7(a) shows that the networks only with ReLU activation function is not able to learn the solution with smooth first-order derivative. However, in Figure 7(b), our method learns a solution with considerable smooth first-order derivative. On the other hand, for the neural network with squared ReLU activation functions, we observed phenomena of gradient explosion and gradient vanishing during training, attributed to the depth of networks and the square operation. Figure 7(c) shows that the network using squared ReLU suffers from gradient vanishing and therefore fails to learn in the training. These comparisons demonstrate that for deep fully connected networks, our DSRN method outperforms networks using only ReLU or squared ReLU. We further visualize the second-order derivative of the DSRN outputs in Figure 7(d). Figure 7 shows that DSRN is able to approximate the smooth function with the square of ReLU activation function only in the last several layers in the neural network. In Table 1, we list the MSE between the neural network solutions and the exact solution: DSRN achieves more accurate approximation than neural networks with only ReLU or square of ReLU. Note that the neural network with ReLU has zero-valued second-order derivative and thus a larger second-order derivative error.
@@ 7.3 Two-dimensional Diffusion equation
(82) $\frac{\partial y}{\partial t}=\frac{\partial^{2}y}{\partial x^{2}}-e^{-t}(\sin(\pi x)-\pi^{2}\sin(\pi x)),\quad x\in[-1,1],\quad t\in[0,1],$
with initial and boundary conditions $y(x,0)=\sin(\pi x),y(-1,t)=y(1,t)=0$. The exact solution is $y(x,t)=e^{-t}\sin(\pi x)$. We use a 12-layer fully-connected network $\mathcal{NN}_{\boldsymbol{\theta}}(x,t)$ to learn the solution $y(x,t)$, with the activation function of the last four hidden layers being the square of ReLU, and all other activation functions being ReLU. Similarly, the loss function is
@@
where $\left\{(x_{i},t_{i},y_{i}),i=1,\dots,N|y_{i}=e^{-t_{i}}\sin(\pi x_{i})\right\}$ are training data. In the training, we use the Adam optimizer with learning step $0.001$. In Figure 8, we plot surface and contours of the DSRN output and the exact solution, respectively.
(a) DSRN output surface
(b) DSRN output contour
(c) Exact solution surface
(d) Exact solution contour
@@
For better comparison, we plot the contours of the DSRN output and the exact solution in one figure, see Figure 9. These comparisons demonstrate that DSRN accurately approximates the smooth solution.
@@ 8 Conclusion
acknowledgement
The work of Y. Y. are supported by National Institute of General Medical Sciences through grant 1R35GM146894. The work of H. Y. was partially supported by the US National Science Foundation under award DMS-2244988, DMS-2206333, and the Office of Naval Research Award N00014-23-1-2007. The work of Y. X. was supported by the Project of Hetao Shenzhen-HKUST Innovation Cooperation Zone HZQB-KCZYB-2020083.

### v4

@@ 7.1 General function approximation
To highlight the advantage of Squared ReLU activation in DSRN for approximation in Sobolev norm, we compare it against a standard ReLU neural network in learning the cubic function $f(x)=x^{3}$ as well as its first- and second-order derivatives. We separately train a 10-layer fully-connected ReLU neural network, and a DSRN, where the last three layers use squared ReLU activations, to approximate the function $f(x)$. Fig. 7 presents the functions learned by the ReLU network (Fig. 7(a)) and DSRN (Fig. 7(d)), the first-order derivative $f^{\prime}(x)=3x^{2}$ estimated by the ReLU neural network (Fig. 7(b)) and DSRN (Fig. 7(e)), and the second-order derivative $f^{\prime\prime}(x)=6x$ estimated by the ReLU neural network (Fig. 7(c)) and DSRN (Fig. 7(f)). The comparison results demonstrate that the DSRN successfully approximates both the cubic function and its derivatives, capturing smoothness in higher-order Sobolev norms. In contrast, the ReLU neural network can only approximate the function up to the first-order derivative, but entirely fails to approximate the second-order derivative due to the vanishing curvature of ReLU. Consequently, ReLU incurs significant error in higher-order Sobolev approximation, whereas DSRN achieves accurate approximation in higher-order Sobolev norm.
((a)) ReLU: $f(x)=x^{3}$
((b)) ReLU: $f^{\prime}(x)=3x^{2}$
((c)) ReLU: $f^{\prime\prime}(x)=6x$
((d)) DSRN: $f(x)=x^{3}$
((e)) DSRN: $f^{\prime}(x)=3x^{2}$
((f)) DSRN: $f^{\prime\prime}(x)=6x$
@@ 7.2 One-dimensional Poisson equation
$\Delta u(x)=-\pi^{2}\sin(\pi x),\quad x\in\left[-1,1\right],$ (7.1)
with boundary conditions u(1)=u(1)=0u(-1)=u(1)=0. The exact solution is $u(x)=\sin(\pi x)$. To be specific, we approximate the solution $u(x)$ by the neural network $\mathcal{NN}_{\bm{\theta}}(x)$ and train the neural network parameters $\bm{\theta}$ to minimize the mean squared error(MSE)
$\frac{1}{N}\sum_{i=1}^{N}\left[\left(\frac{d^{2}}{dx^{2}}\mathcal{NN}_{\bm{\theta}}(x_{i})+\pi^{2}\sin(\pi x_{i})\right)^{2}+\left(\mathcal{NN}_{\bm{\theta}}(x_{i})-y_{i}\right)^{2}\right].$
@@
((a)) ReLU
((b)) DSRN
((c)) Squared ReLU
((d)) DSRN
@@
In Fig. 8, We plot the prediction of the neural network with ReLU (Fig. 8(a)), DSRN (Fig. 8(b)) and the square of ReLU (Fig. 8(c)), respectively. We also plot the first-order derivative $\frac{d}{dx}\mathcal{NN}_{\bm{\theta}}$ using the autograd. The exact solution is $u(x)=\sin(\pi x)$, which corresponds to the ground truth in the figure. Fig. 8(a) shows that the networks only with ReLU activation function is not able to learn the solution with smooth first-order derivative. However, in Fig. 8(b), our method learns a solution with considerable smooth first-order derivative. On the other hand, for the neural network with squared ReLU activation functions, we observed phenomena of gradient explosion and gradient vanishing during training, attributed to the depth of networks and the square operation. Fig. 8(c) shows that the network using squared ReLU suffers from gradient vanishing and therefore fails to learn in the training. These comparisons demonstrate that for deep fully connected networks, our DSRN method outperforms networks using only ReLU or squared ReLU. We further visualize the second-order derivative of the DSRN outputs in Fig. 8(d). Fig. 8 shows that DSRN is able to approximate the smooth function with the square of ReLU activation function only in the last several layers in the neural network. In Table 2, we list the MSE between the neural network solutions and the exact solution: DSRN achieves more accurate approximation than neural networks with only ReLU or square of ReLU. Note that the neural network with ReLU has zero-valued second-order derivative and thus a larger second-order derivative error.
@@ 7.3 Two-dimensional Diffusion equation
$\frac{\partial y}{\partial t}=\frac{\partial^{2}y}{\partial x^{2}}-e^{-t}(\sin(\pi x)-\pi^{2}\sin(\pi x)),\quad x\in[-1,1],\quad t\in[0,1],$ (7.2)
with initial and boundary conditions $y(x,0)=\sin(\pi x),y(-1,t)=y(1,t)=0$. The exact solution is $y(x,t)=e^{-t}\sin(\pi x)$. We use a 12-layer fully-connected network $\mathcal{NN}_{\bm{\theta}}(x,t)$ to learn the solution $y(x,t)$, with the activation function of the last four hidden layers being the square of ReLU, and all other activation functions being ReLU. Similarly, the loss function is
@@
where $\left\{(x_{i},t_{i},y_{i}),i=1,\dots,N|y_{i}=e^{-t_{i}}\sin(\pi x_{i})\right\}$ are training data. In the training, we use the Adam optimizer with learning step $0.001$. In Fig. 9, we plot surface and contours of the DSRN output and the exact solution, respectively.
((a)) DSRN output surface
((b)) DSRN output contour
((c)) Exact solution surface
((d)) Exact solution contour
@@
For better comparison, we plot the contours of the DSRN output and the exact solution in one figure, see Fig. 10. These comparisons demonstrate that DSRN accurately approximates the smooth solution.
@@ 8 Conclusion
Acknowledgments
The work of H. Y. was partially supported by the US National Science Foundation under award DMS-2244988, DMS-2206333, and the Office of Naval Research Award N00014-23-1-2007. The work of Y. X. was supported by the Project of Hetao Shenzhen-HKUST Innovation Cooperation Zone HZQB-KCZYB-2020083.